OpenStudy (anonymous):

help me Find the median for the given set of data. 0.6, 1.1, 0.87, 2, 0.87, 1.23

OpenStudy (anonymous):

add them up then divide by number of terms

OpenStudy (anonymous):

thats the mean

OpenStudy (anonymous):

0.6+1.1+0.87+2+0.87+1.23= 6.67/6=1.111666666666 repeating. so 1.1 basically

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

thats the mean

OpenStudy (texaschic101):

median is the middle number...so how many total numbers do you have ?

OpenStudy (anonymous):

apparently it is \(.985\) http://www.wolframalpha.com/input/?i=median+0.6%2C+1.1%2C+0.87%2C+2%2C+0.87%2C+1.23

OpenStudy (anonymous):

6

OpenStudy (anonymous):

oops sorry guys I did the mean instead of the median. well for the median order them from least to greatest

OpenStudy (texaschic101):

First put the numbers in order from smallest to largest.. if there is an even number of numbers, then the median can be found by taking the 2 middle numbers and dividing by 2. Can you do that ?

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

really ? why

OpenStudy (anonymous):

my math skills are not that good that is why i use wolfram

OpenStudy (anonymous):

0.6,0.87,0.87,1.1,1.23,2 what I do is cross out one from each side till u get to the middle which is the same thing as finding the middle number in this case there r 2 middle numbers. 0.87 and 1.1 add them together then divide by 2 0.87+1.1= 1.97/2= 0.985 I think that's the answer

OpenStudy (texaschic101):

I agree.....just like I said...add the 2 middle numbers and divide by 2

OpenStudy (texaschic101):

now if there was an odd number of numbers, there will be only 1 middle number, and that number would be the median